OpenStudy (anonymous):

solve 5/(5-p) - p^2/(5-p) = -2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not one

OpenStudy (anonymous):

just a min i will solve it

OpenStudy (nurali):

\[\frac{ 5 }{ 5-p }-\frac{ p^2 }{ 5-p }=-2\] \[\frac{ 5-p^2 }{ 5-p }=-2\] \[5-p^2=-2(5-p)\] \[5-p^2=-10+2p\] \[p^2+2p-15=0\] \[p^2+5p-3p-15=0\] \[p(p+5)-3(p+5)=0\] \[(p+5)(p-3)=0\] \[p=-\]

OpenStudy (nurali):

\[p=-5,3\]

OpenStudy (skullpatrol):

@Nurali @blkmrc Do you REALLY think that doing all the work will help someone learn?

OpenStudy (nurali):

how 5-p^2=(5-p)(5+p)

OpenStudy (nurali):

(5-p)(5+p)=25-p^2

OpenStudy (anonymous):

@skullpatrol it really does help because I can see how they get the answer

OpenStudy (anonymous):

ohhhhhhhhhhh you have righ i am sorry everybody I have made mistake :/

OpenStudy (skullpatrol):

@blkmrc keep HOPING

OpenStudy (anonymous):

I just woke up I am sorry for my answer shame for me thanks for warning @skullpatrol @Nurali

OpenStudy (skullpatrol):

Open Study Code of Conduct Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (skullpatrol):

@Alexis2332 During a test you can look over at someone else's work and see how to get the answer to, right?

OpenStudy (skullpatrol):

$$learning \neq copying$$

OpenStudy (e.cociuba):

Very true skull. I encountered one of these problems yet again today.